# NeuraLeaf: Neural Parametric Leaf Models with Shape and Deformation Disentanglement

### v1

@@ 1 Introduction
To these issues, we integrate leaves’ bio-inspired insights with a data-driven framework. We propose a leaf-specific NPM representation, NeuraLeaf, leveraging the nature of leaves, i.e., the flattened leaves can be approximated as a plane. As illustrated in Fig. 1, we disentangle the leaves’ 3D geometry into their 2D base (i.e., flattened) shape and 3D deformation represented by different latent codes.
@@ 2.1 Parametric Models
Blanz et al. [4] propose an early attempt at a parametric model representing facial variations using a low-dimensional shape space extracted by principal component analysis (PCA). Along with the improvements in the 3D registration technique [46] and large-scale data capture [1, 49], parametric 3D models have emerged as a leading method to model domain-specific deformable 3D shapes. The parametric models are studied in wide fields like human bodies [1, 23], hands [40], animals [59], faces [22], and leaves [5, 3]. SMPL [23], a representative example of parametric human body models, combines a low-dimensional shape space with an articulated blend-skinned model and is used in various applications [38, 25, 47]. However, such skinned vertex-based models have limited resolution and have difficulty representing non-rigid surface deformations often seen in leaves.
@@ 2.2 Leaf Modeling
Traditionally, B-spline curves were used for representing silhouettes or veins of 2D leaves [19, 27], while they require tailor-made parameter sets for each group of leaves. For leaf modeling applicable to various leaves, Barley et al. [5] propose a parametric leaf model using a traditional PCA-based model [4] for leaf variation within a given flora species. Gaurav et al. [3] extend this direction for representation of non-planar leaves by using the Bezier curve, emphasizing identifying three pivotal features of leaf geometry: midrib, left/right silhouette, and the cross-section. However, PCA-based models are constrained by limited shape variability, lack deformation modeling, and require large 3D datasets. While offering parametric control of deformation, Bezier curve-based models still cannot capture the complex and natural deformations of real leaves. Compared to existing works, our work is the first to propose NPMs for leaves, representing natural and diverse shapes and deformation.
@@
Existing methods for leaf deformation modeling often derive from physically-based simulations based on mesh. For instance, venation skeletons and surface membranes are used to model aging leaves [14], allowing for interactive user-driven deformation. Similarly, [24] introduces a technique to animate leaf wilting due to gravity, also employing a venation skeleton for deformation. Jeong et al. [45] generate intricate wrinkles and folds in leaves by simulating inhomogeneous shrinkage, which is influenced by physical properties like loss of water. Beyond those limited variations of deformations, our approach learns an implicit representation of deformation directly from real scanned leaf data, enabling high-fidelity and general-purpose modeling and reconstruction.
@@ 3.3 Deformation Space
Figure 3: Deformation mapping in the zero-mean latent space. Samples closer to the center exhibit smaller deformations, while those farther from the center show greater deformations.
Besides common constraints, we newly propose an additional regularization term, deformation mapping loss $\mathcal{L}_{map}$, to align the deformation latent space. Inspired by the fact that the deformation space forms a zero-mean Gaussian, we measure the chamfer distance $d_{cham}$ between the base and deformed shape $(\mathcal{S}_{b},\mathcal{S}_{d})$. The loss function maps the amount of deformation to the Gaussian distribution as
@@
Second stage: The second stage aims to generalize the learned deformation space to the entire base shape space, including those without explicitly paired deformation samples. Due to the significant differences between base shapes, minimizing the Chamfer distance between dissimilar leaf pairs without altering the base shape is challenging. We thus select $m$ similar base shapes from the entire base shape space for each base shape in DeformLeaf based on the IoU between base masks, and we encourage similar base shapes to have similar skinning weights during training. We find that using only chamfer distance causes the unpaired base shape to overly conform to the deformed shape. We, therefore, introduce two additional boundary-focused constraints, boundary length loss $\mathcal{L}_{\text{bound}}$ and face angle loss $\mathcal{L}_{\text{ang}}$ to maintain smoothness and preserve base shapes.
@@
$\vspace{-2mm}\mathcal{L}_{\text{bound}}=\sum_{\mathbf{v}^{\prime}\in\mathcal{V}^{\prime}}\left(\|\mathbf{v}_{i}^{\prime}-\mathbf{v}^{\prime}_{i+1}\|_{2}-\|\mathbf{v}_{i}-\mathbf{v}_{i+1}\|_{2}\right)^{2},$ (12)
@@ 4 DeformLeaf Dataset
Figure 4: Pipeline of dataset acquisition.
@@
Our DeformLeaf dataset comprises approximately $300$ base-deformation leaf pairs, the capture setting is shown in Fig. 4. Multi-view images of a leaf are taken from various angles to gather comprehensive visual data, and a shot from a top-down view of the base image is also taken with a light board. The 3D reconstruction process is performed by Metashape [^1] Agisoft Metashape, https://www.agisoft.com/, last accessed on March 5, 2025..
@@ 5 NeuraLeaf Fitting to 3D Observations
Aside from pure CG modeling, NeuraLeaf can be used for reconstruction tasks by fitting our model to real-world observations. Especially in plant science and agriculture, beyond ordinary 3D shape reconstruction such as by multi-view stereo or RGB-D imaging, it is crucial to acquire 3D structural information such as topology, number, area, relationship, and direction of leaves from reconstructed shape information, in which our model helps.
We, therefore, provide a model fitting pipeline of NeuraLeaf to 3D shapes. The goal of NPM fitting is to find the set of latent codes, $(\mathbf{z}_{s},\mathbf{z}_{d})$ in our case, fit to the observations. To this end, our pipeline consists of the neural-net-based inversion of latent codes, followed by a direct optimization of the latent codes for refinement. We also take a similar fitting strategy for the texture latent $\mathbf{z}_{t}$ but without the latent code inversion, which is detailed in the supplementary materials.
@@ 6.2 Results
Method $C$-$\ell_{2}$ [mm] $\downarrow$ NC $\uparrow$ Corres-free Temp-free Inf. time [s]
@@
To mimic practical use, we perform leaf reconstruction from observation of multiple leaves, which contain occlusions of leaves. We use top-view RGB-D observations of $5$ plants, where we provide segmentation masks and obtain point clouds corresponding to each leaf instance. For the ground-truth shapes, we reconstruct the plants’ shapes by multi-view stereo (MVS) using many ($\geq 100$) images captured under a windless environment and to accurately recover the occlusions occurring in the top-view images. The fitting process for every single leaf employs the same workflow described in Sec. 5, while enhancing to share similar shape latent $\mathbf{z}_{s}$ for the observation of the same plant species, whose details are described in the supplementary materials. For the NPM baseline, we also share the shape latent in a similar manner to our method.
@@
Figure 6: Visual comparisons of multiple leaf reconstruction. We use top-view RGB-D observations as input. Carefully captured multiview images are used to create the ground truths (denoted as “dense point cloud”). Our method achieves faithful reconstruction (e.g., vein shapes on the leaf surfaces) and better occlusion recovery (e.g., middle leaf in the bottom row).
@@
Method $C$-$\ell_{2}$ [mm] $\downarrow$ NC $\uparrow$
@@
Figure 7: Deformation transfer results. Each row in (a) shows different leaf species with the same deformation latent. (b) demonstrate that each deformed leaf shape can be represented by a general surface, supporting the idea of a generalized deformation space capable of all kinds of leaf shapes.
@@ 6.3 Ablation Study
We perform an ablation study to assess the effect of the deformation mapping loss $\mathcal{L}_{\text{map}}$ used for the training of deformation space. Table 3 (top) shows the results of the ablation studies and demonstrates that deformation mapping loss contributes to the improvement of shape modeling.
@@
Method $C$-$\ell_{2}$ [mm] $\downarrow$ NC $\uparrow$
@@
# of control points $C$-$\ell_{2}$ [mm] $\downarrow$ NC $\uparrow$
@@ A.3 Non-rigid Registration
As mentioned in Sec. 4 in the main paper, we perform a non-rigid registration pipeline using the As-rigid-as-possible (ARAP) registration followed by the Coherent Point Drift (CPD) registration. Since the leaves have flexible deformation, it is often challenging to yield a reasonable registration by a single method. We, therefore, combine two registration methods to gradually relax the degrees of freedom in the deformation, namely, rigid alignment, ARAP assuming a locally rigid shape, and CPD as non-rigid registration.
@@
where $\hat{\textbf{v}}_{i}=\textbf{v}_{i}+\delta_{i}$, $d(\hat{\textbf{v}}_{i}-\mathbf{v}^{\prime}_{i})$ is the mean point-to-point distance from $\hat{\textbf{v}}_{i}$ to its nearest neighbors in $\mathcal{V}^{\prime}$, and $\mathcal{N}_{\mathbf{v}_{i}}$ denotes the nearest neighbouring vertices to ${\mathbf{v}_{i}}$.
@@ B.1 Architecture of Decoders
Shape decoder $f_{\theta_{s}}$ is an MLP that consists of $8$ layers, with a dimension of hidden layers of $512$ and a single skip connection from the input to the middle layer, which is similar to an existing NPM [8]. The initialization strategy of parameters of $f_{\theta_{s}}$ is the same as [2] so that $f_{\theta_{s}}$ produces an approximate SDF of a unit sphere.
@@ C.4 Details for Multiple Leaf Reconstruction
To begin the reconstruction process, each leaf instance is initially processed by the shape encoder $f_{\Omega_{s}}$ to estimate the shape latent code $\mathbf{z}_{s}$. However, due to occlusion or overlapping regions among leaves, some of the estimated shape codes might not accurately represent the true leaf shape. To identify a representative anchor latent code in the presence of potential outliers, we first compute the similarity between latent codes using Euclidean distance and perform clustering. Specifically, we employ K-means clustering to divide the latent codes into $K=3$ clusters and select the cluster with the largest number of latent codes as the main cluster, assuming that this cluster represents the normal shape. Once the main cluster is identified, we choose an anchor latent code $\mathbf{z}_{s}^{anc}$, by selecting the code closest to the cluster center. $\mathbf{z}_{s}^{anc}$ serves as the reference during optimization. Formally, we introduce a regularization term $\mathcal{L}_{anc}$ defined as
@@
Method $C$-$\ell_{2}$ [mm] $\downarrow$ NC $\uparrow$

### v2

@@ 1 Introduction
To address them, we integrate leaves’ biological insights with a data-driven framework. We propose a leaf-specific NPM representation, NeuraLeaf, leveraging the nature of leaves, i.e., the flattened leaves can be approximated as a plane. As illustrated in Fig. 1, we disentangle the leaves’ 3D geometry into their 2D base (i.e., flattened) shape and 3D deformation represented by different latent codes.
@@ 2.1 Parametric Models
Blanz et al[4] propose an early attempt at a parametric model representing facial variations using a low-dimensional shape space extracted by principal component analysis (PCA). Along with the improvements in the 3D registration technique [46] and large-scale data capture [1, 49], parametric 3D models have emerged as a leading method to model domain-specific deformable 3D shapes. The parametric models are studied in wide fields like human bodies [1, 23], hands [40], animals [59], faces [22], and leaves [5, 3]. SMPL [23], a representative example of parametric human body models, combines a low-dimensional shape space with an articulated blend-skinned model and is used in various applications [38, 25, 47]. However, such skinned vertex-based models have limited resolution and have difficulty representing non-rigid surface deformations often seen in leaves.
@@ 2.2 Leaf Modeling
Traditionally, B-spline curves were used for representing silhouettes or veins of 2D leaves [19, 27], while they require tailor-made parameter sets for each group of leaves. For leaf modeling applicable to various leaves, Barley et al[5] propose a parametric leaf model using a traditional PCA-based model [4] for leaf variation within a given flora species. Gaurav et al[3] extend this direction for the representation of non-planar leaves by using the Bezier curve, emphasizing identifying three pivotal features of leaf geometry: midrib, left/right silhouette, and the cross-section. However, PCA-based models are constrained by limited shape variability, lack deformation modeling, and require large 3D datasets. While offering parametric control of deformation, Bezier curve-based models still cannot capture the complex and natural deformations of real leaves. Compared to existing works, our work is the first to propose NPMs for leaves, representing natural and diverse shapes and deformations.
@@
Existing methods for leaf deformation modeling derive from physically-based simulations. For instance, venation skeletons and surface membranes are used to model aging leaves [14], allowing for interactive user-driven deformation. Similarly, [24] animates leaf wilting due to gravity using a venation skeleton for deformation. Jeong et al[45] generate intricate wrinkles and folds in leaves by simulating inhomogeneous shrinkage, which is influenced by physical properties like loss of water. Beyond those limited variations of deformations, our approach learns an implicit representation of deformation directly from real scanned leaf data, enabling high-fidelity and general-purpose modeling and reconstruction.
@@ 3.3 Deformation Space
Figure 3: Deformation mapping in the zero-mean latent space. The colormap indicates the Chamfer distance between each deformed shape and its corresponding base shape. Samples closer to the center exhibit smaller deformations, while those farther from the center show greater deformations.
Besides common constraints, we newly propose an additional regularization term, deformation mapping loss $\mathcal{L}_{map}$, to align the deformation latent space. Inspired by the fact that the deformation space forms a zero-mean Gaussian, we measure the chamfer distance $d_{cham}$ between the base and deformed shape $(\mathcal{S}_{b},\mathcal{S}_{d})$. This loss maps the amount of deformation to the Gaussian distribution as
@@
Second stage: This stage aims to generalize the learned deformation space to the entire base shape space, including those without explicitly paired deformation samples. Due to the significant differences between base shapes, minimizing the Chamfer distance between dissimilar leaf pairs without altering the base shape is challenging. We thus select $m$ similar base shapes from the entire base shape space for each base shape in DeformLeaf based on the IoU between base masks, and we encourage similar base shapes to have similar skinning weights during training. We find that using only the Chamfer distance causes the unpaired base shape to overly conform to the deformed shape. We, therefore, introduce two additional boundary-focused constraints, boundary length loss $\mathcal{L}_{\text{bound}}$ and face angle loss $\mathcal{L}_{\text{ang}}$ to maintain smoothness and preserve base shapes.
@@
$\vskip-5.69054pt\mathcal{L}_{\text{bound}}=\sum_{\mathbf{v}^{\prime}\in\mathcal{V}^{\prime}}\left(\|\mathbf{v}_{i}^{\prime}-\mathbf{v}^{\prime}_{i+1}\|_{2}-\|\mathbf{v}_{i}-\mathbf{v}_{i+1}\|_{2}\right)^{2},$ (12)
@@ 4 DeformLeaf Dataset
Figure 4: Capturing pipeline for paired base and deformed shapes.
@@
Our DeformLeaf dataset comprises approximately $300$ base-deformation leaf pairs; the capture setting is shown in Fig. 4. Multi-view images of a leaf are taken from various angles to gather comprehensive visual data, and a shot from a top-down view of the base image is also taken with a light board. The 3D reconstruction process is performed by Metashape [^1] Agisoft Metashape, https://www.agisoft.com/, last accessed on March 5, 2025..
@@ 5 NeuraLeaf Fitting to 3D Observations
Aside from pure CG modeling, NeuraLeaf can be used for reconstruction tasks by fitting our model to real-world observations. Especially in plant science and agriculture, beyond ordinary 3D shape reconstruction, such as by multi-view stereo or RGB-D imaging, it is crucial to acquire 3D structural information, such as topology, number, area, relationship, and direction of leaves from reconstructed shape information, in which our model helps.
The goal of NPM fitting is to find the set of latent codes, $(\mathbf{z}_{s},\mathbf{z}_{d})$ in our case, fit to the observations. To this end, our pipeline consists of the neural-net-based inversion of latent codes, followed by a direct optimization of the latent codes for refinement. We also take a similar fitting strategy for the texture latent $\mathbf{z}_{t}$ but without the latent code inversion, which is detailed in the supplementary materials.
@@ 6.2 Results
Method $C$-$\ell_{2}$ [$m$] $\downarrow$ NC $\uparrow$ Corres-free Temp-free Inf. time [s]
@@
To mimic practical use, we perform leaf reconstruction from observation of multiple leaves, which contain occlusions of leaves. We use top-view RGB-D observations of $5$ plants, where we provide segmentation masks and obtain point clouds corresponding to each leaf instance. For the ground-truth shapes, we reconstruct the plants’ shapes by multi-view stereo (MVS) using many ($\geq 100$) images captured under a windless environment to accurately recover the occlusions occurring in the top-view images. The fitting process for every single leaf employs the same workflow described in Sec. 5, while enhancing to share similar shape latent $\mathbf{z}_{s}$ for the observation of the same plant species, whose details are described in the supplementary materials. For the NPM baseline, we also share the shape latent in a similar manner to our method.
@@
Figure 6: Visual comparisons of multiple leaf reconstruction. We use top-view RGB-D observations as input. Carefully captured multi-view images are used to create the ground truths (denoted as “dense point cloud”). Our method achieves faithful reconstruction (e.g., vein shapes on the leaf surfaces) and better occlusion recovery (e.g., middle leaf in the bottom row).
@@
Method $C$-$\ell_{2}$ [$m$] $\downarrow$ NC $\uparrow$
@@
Figure 7: Deformation transfer results. Each row in (a) shows different leaf species with the same deformation latent. (b) demonstrates that each deformed leaf shape in a different color can be represented by a general surface, supporting the idea of a generalized deformation space capable of all kinds of leaf shapes.
@@ 6.3 Ablation Study
We perform an ablation study to assess the effect of the deformation mapping loss $\mathcal{L}_{\text{map}}$ used for the training of the deformation space. Table 3 (top) shows the results of the ablation studies and demonstrates that deformation mapping loss contributes to the improvement of shape modeling.
@@
Method $C$-$\ell_{2}$ [$m$] $\downarrow$ NC $\uparrow$
@@
# of control points $C$-$\ell_{2}$ [$m$] $\downarrow$ NC $\uparrow$
@@ A.3 Non-rigid Registration
As mentioned in Sec. 4 in the main paper, we perform a non-rigid registration pipeline using the As-rigid-as-possible (ARAP) registration followed by the Coherent Point Drift (CPD) registration. Since the leaves have flexible deformation, yielding a reasonable registration by a single method is often challenging. Therefore, we combine two registration methods to relax the degrees of freedom in the deformation gradually: rigid alignment, ARAP assuming a locally rigid shape, and CPD as non-rigid registration.
@@
where $\hat{\textbf{v}}_{i}=\textbf{v}_{i}+\delta_{i}$, $d(\hat{\textbf{v}}_{i}-\mathbf{v}^{\prime}_{i})$ is the mean point-to-point distance from $\hat{\textbf{v}}_{i}$ to its nearest neighbors in $\mathcal{V}^{\prime}$, and $\mathcal{N}_{\mathbf{v}_{i}}$ denotes the nearest neighboring vertices to $\mathbf{v}_{i}$}.
@@ B.1 Architecture of Decoders
Shape decoder $f_{\theta_{s}}$ is an MLP that consists of $8$ layers, with a dimension of hidden layers of $512$ and a single skip connection from the input to the middle layer, which is similar to an existing NPM [8]. The initialization strategy of parameters of $f_{\theta_{s}}$ is the same as [2], so that $f_{\theta_{s}}$ produces an approximate SDF of a unit sphere.
@@ C.4 Details for Multiple Leaf Reconstruction
To begin the reconstruction process, each leaf instance is initially processed by the shape encoder $f_{\Omega_{s}}$ to estimate the shape latent code $\mathbf{z}_{s}$. However, due to occlusion or overlapping regions among leaves, some of the estimated shape codes might not accurately represent the true leaf shape. To identify a representative anchor latent code in the presence of potential outliers, we first compute the similarity between latent codes using Euclidean distance and perform clustering. Specifically, we employ K-means clustering to divide the latent codes into $K=3$ clusters and select the cluster with the largest number of latent codes as the main cluster, assuming that this cluster represents the normal shape. Once the main cluster is identified, we choose an anchor latent code $\mathbf{z}_{s}^{anc}$ by selecting the code closest to the cluster center. $\mathbf{z}_{s}^{anc}$ serves as the reference during optimization. Formally, we introduce a regularization term $\mathcal{L}_{anc}$ defined as
@@
Method $C$-$\ell_{2}$ [$m$] $\downarrow$ NC $\uparrow$